### v2

# Penalizing Length: Uncovering Systematic Bias in Quality Estimation Metrics

Abstract
Quality Estimation (QE) metrics are vital in machine translation for reference-free evaluation and as a reward signal in tasks like reinforcement learning. However, the prevalence and impact of length bias in QE have been underexplored. Through a systematic study of top-performing regression-based and LLM-as-a-Judge QE metrics across 10 diverse language pairs, we reveal two critical length biases: First, QE metrics consistently over-predict errors with increasing translation length, even for high-quality, error-free texts. Second, they exhibit a preference for shorter translations when multiple candidates are available for the same source text. These inherent length biases risk unfairly penalizing longer, correct translations and can lead to sub-optimal decision-making in applications such as QE reranking and QE guided reinforcement learning. We further investigate the root cause, presenting evidence that QE models conflate quality with sequence length due to skewed supervision distributions. As a diagnostic intervention, we apply length normalization during training. We show that this simple intervention is sufficient to decouple error probability from length, effectively counteracting the data skew and yielding more reliable QE signals.
Machine Translation (MT), Quality Estimation, LLM-as-a-Judge
Figure 1: QE metrics exhibit a length-related penalty when evaluating long-form translations, as metric scores decrease when translations become longer despite being error-free. As shown, concatenating consecutive gold-standard segments from the same document leads to progressively lower predicted scores, in contrast to human evaluations.
1 Introduction
Quality Estimation (QE) predicts the quality of machine-translated text without relying on reference translations, serving as a vital signal for applications where ground truth is unavailable. Broadly, existing QE methods fall into two categories: regression-based models, such as MetricX (Juraska et al., 2024) and COMET (Rei et al., 2020), which learn to assess the hypothesis directly from annotated data, and LLM-as-a-Judge approaches, such as InstructScore (Xu et al., 2023) and AutoMQM (Fernandes et al., 2023), which leverage large language models to generate error ratings that approximate human judgments.
Length bias refers to the systematic tendency of evaluation models to prefer or penalize outputs depending on their length rather than their true quality. Prior research has documented such effects in related evaluation settings, including LLM-as-a-Judge frameworks (Domhan and Zhu, 2025) and reward models used in reinforcement learning (Zhao et al., 2025). However, the extent to which QE metrics suffer from length bias remains largely underexplored. This gap is consequential, as length bias can distort system comparisons, over- or under-value longer translations, and ultimately misguide both model development and data curation.
In this study, we design two complementary experiments to systematically investigate the presence of length bias in QE metrics. The first experiment evaluates length invariance by concatenating high-quality, error-free segments into progressively longer translations, allowing us to assess whether metrics remain stable as input length increases. The second experiment directly compares alternative hypotheses generated from the same source, enabling us to measure whether metrics exhibit a consistent preference toward shorter or longer outputs. Our analyses reveal two concerning patterns: QE metrics tend to assign spurious errors as translations grow longer, even when no actual errors are present, and they disproportionately favor shorter hypotheses when multiple candidates exist. These findings highlight a fundamental vulnerability: QE metrics exhibit a systematic length penalty, assigning spurious errors to longer segments and disproportionately favoring shorter hypotheses in reranking scenarios.
To address these shortcomings, we analyze the training data of a representative state-of-the-art QE metric, MetricX-24, and identify an implicit label imbalance. Specifically, supervision signals are unevenly distributed across sequence lengths, inducing a spurious association whereby longer translations are more likely to contain errors and thus receive worse labels. Motivated by this diagnosis, we apply length normalization during training to counteract this length-dependent skew. Experiments show that length-normalized training substantially reduces length sensitivity and yields more reliable QE signals across translations of varying lengths.
2 Related Work
2.1 Length Effects in MT Evaluation Metrics
Early reference-based MT metrics have long been sensitive to translation length. For instance, BLEU introduces a brevity penalty (BP) to discourage overly short hypotheses that could artificially inflate n-gram precision (Papineni et al., 2002). Likewise, chrF uses an F-score over character-n-gram precision and recall, penalizing outputs that are too short (low recall) or too long (lower precision via extra content or mismatches) relative to the reference (Popović, 2015). Nonetheless, length remains a difficult dimension to handle. (Papineni et al., 2002) note that neither the brevity penalty nor modified precision alone fully enforces appropriate length matching. Moreover, recent work at the document level shows that BLEU can mask quality improvements when output length increases: for example, (Sun et al., 2025) find that translating entire documents produced better human-judged quality, yet BLEU scores often did not reflect this improvement, and sometimes ranked shorter, less coherent translations higher.
Beyond reference-based metrics, learned neural metrics and LLM-based evaluators also display length sensitivity. For example, Karpinska et al. (2022) reveal that even state-of-the-art metrics such as COMET can struggle with subtle perturbations, including omissions or repetitions, which may occur more frequently in unusually short or long translations. For document-level translations, (Domhan and Zhu, 2025) show that when using LLM-as-a-Judge evaluators, longer input texts tend to result in fewer detected error spans and lower system-ranking accuracy, indicating a bias related to text length rather than translation quality.
These findings underscore that handling translation length properly remains an important and under-addressed challenge in MT evaluation. In our work, we therefore situate ourselves within this context by investigating whether current QE methods penalize translations simply for being longer (or shorter) rather than being inherently worse.
2.2 Biases in Quality Estimation Models
Modern QE models have evolved from feature-based frameworks such as QuEst (Specia et al., 2013) to neural architectures that achieve strong results on recent benchmarks (Rei et al., 2022, 2023), with representative examples including OpenKiwi, TransQuest, and CometKiwi (Kepler et al., 2019; Ranasinghe et al., 2020; Rei et al., 2022). Despite these advances, numerous studies have noted that learned QE models often pick up spurious correlations or biases from their training data. For instance, (Behnke et al., 2022) identify a partial-input bias: when trained on standard data, QE models may over-rely on features of the target translation (e.g., grammaticality) rather than verifying adequacy or fidelity to the source, leading them to assign high quality scores to fluent outputs that fail to preserve meaning.
Such biases often stem from limitations in the training data. QE models are typically supervised on human ratings of MT outputs, such as WMT Direct Assessments or MQM annotations. When these datasets are skewed or unrepresentative, the resulting QE models tend to mirror those biases. For instance, recent work finds that automatic metrics frequently assign lower scores to human-generated translations than expected, relative to MT outputs, implying a bias against human style (Deutsch et al., 2025). This bias arises because the metrics are predominantly trained on machine-generated outputs and thus fail to capture the nuances of human translations, which often display greater lexical and structural diversity. Similarly, domain and language biases have been reported: metrics can overfit to the score distributions of particular domains or language pairs (Zouhar et al., 2024), limiting their generalization to new settings without adaptation. Complementary to these findings, we examine whether the observed length bias in QE correlates with the under-representation of long-form, error-free examples in the training data.
3 Preliminaries
Definition of Length Bias.
Length bias describes a systematic preference for certain output lengths irrespective of true quality. Formally, we evaluate the length bias of QE models using the statistical bias definition. This bias is characterized by the disparity between a predicted quality score and a true quality score, as follows:
$\textrm{Bias}(\hat{\theta})=\mathbb{E}[\hat{\theta}]-\theta$ (1)
[…]
4 Experiments
Task Definition.
A QE model, $M_{\text{QE}}$, assesses the quality of a system’s hypothesis $h$ given a source text $x$ without relying on human references, assigning a score based on predefined criteria. In our experimental setup, we investigate the phenomenon of length bias by comparing the scores of groups of source–hypothesis pairs $(x,h)$ with the same human-rated translation quality, and the instances in each group are formed by either varying the sequence length of error-free translations or selecting alternative hypotheses of differing lengths for the same source.
4.1 Experiment Setup
Dataset.
[…]
Quality Estimation Metrics.
We consider four popular QE metrics spanning two major evaluation frameworks: (i) two regression-based metrics, MetricX-24 QE (Juraska et al., 2024) and CometKiwi (Rei et al., 2022), and (ii) two LLM-as-a-Judge variants, AutoMQM fine-tuned on Gemini-2.0 (Fernandes et al., 2023; Finkelstein et al., 2024) and directly prompted Gemini-2.5-pro model (Gemini Team, 2025), both of which are used to perform MQM-style evaluation.
Hypotheses.
We investigate two forms of length bias in this study: (i) bias with respect to source text length and (ii) bias with respect to translation length. For (i), we construct multiple source–translation pairs of equivalent quality to test whether the metrics show systematic preferences related to source text length. For (ii), we compare pairs of translations of equal quality generated from the same source, categorize them by relative length, and analyze whether the metrics display a consistent preference for a particular category.
4.2 Bias in Source Text Length
Figure 2: Trends of metric scores with increasing source text length across different metrics. Each data point denotes the score change relative to the first passage.
Experiment Setup.
To examine bias in source text length, we construct a dataset from the WMT24++ data as follows: for each language pair, we collect the first 5 segments from each document in the data, denoted as $s_{1}$, $s_{2}$, $s_{3}$, $s_{4}$ and $s_{5}$, where each segment is approximately 1-2 sentences long. Then, we reconstruct the original document flow by gradually appending the subsequent contiguous segment $s_{i}$ to the previous sequence. Formally, for each document $d$, we construct a group of 5 passages, denoted as $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ and $p_{5}$, where $p_{i}=s_{1}+\dots+s_{i}$. This ensures that the resulting passages are not arbitrary concatenations, but represent the original, coherent discourse intended by the source texts.
We construct the dataset as such with two considerations. First, selecting segments from the same document keeps the constructed passages within a single domain and minimizes confounds from topic variation. Second, to test our hypothesis, the dataset must include source–translation pairs spanning a wide range of lengths; thus, we concatenate consecutive segments to preserve inter-segment fluency better and reduce boundary-induced disfluency that could potentially depress performance.
Since the metrics inherit different context-window limits from their base models (e.g., 512 tokens for MetricX-24 QE), we standardize the context window to the smallest across metrics, and thus discard documents whose first 5 segments exceed 500 tokens using MetricX-24’s tokenizer  [^1] https://huggingface.co/google/mt5-xl. We then evaluate each group of 5 passages using distinct metrics and analyze the relationship between source text length and translation score. To control for score variations between language pairs, we normalize scores within each group by subtracting the score of $p_{1}$ from $p2$, $p3$, $p4$, and $p5$, and report the score differences instead of raw scores. Figure 2 presents the trend of performance change with line charts over passage index across language pairs from all metrics.
Figure 3: Trends of MetricX-24 QE score with different severity/error types, evaluated on WMT24++ En-De and En-Zh.
Figure 4: Trends of Gemini AutoMQM score with different severity/error types, evaluated on WMT24++ En-De and En-Zh.
Results.
As shown in Figure 2, we observe a consistent trend across all language pairs: translation quality deteriorates as source text length increases. Specifically, extending the source text from 1 to 5 segments results in score drops of 4.0–8.0 points for MetricX-24 QE, 1.5–5.6 points for CometKiwi, 1.7–5.3 points for Gemini AutoMQM, and 1.8–8.5 points for Gemini LLM-as-a-Judge.
[…]
To better quantify this behavior, we note that the metrics utilize the MQM error scheme (Freitag et al., 2021), which assigns penalties of –1 (minor) and –5 (major) on a 25-point scale; normalized to our 100-point scale, these correspond to penalties of –4 and –20, respectively. While a marginal decrease in quality scores might be expected due to the cumulative probability of minor imperfections in longer sequences ($P^{N}_{error-free}$), the magnitude of the observed drop, particularly in LLM-as-a-Judge metrics (up to 8.5 points), far exceeds the theoretical decay of joint error probability. This suggests the penalty is driven by a systematic bias against length rather than a valid accumulation of errors.
Figure 5: Average preference for shorter translations by length difference.
4.3 Bias in Error Type and Severity
The preceding experiments show that length bias exists relative to source text length. What remains unclear, however, is whether this bias persists when translations contain explicit errors. To address this question, we further investigate whether the bias interacts with specific combinations of error severity and error type. Accordingly, we design a follow-up experiment introducing four controlled categories of errors. Building on the setup described in Section 4.2, we modify the first segment of each passage by inserting one of the following: (a) a major accuracy error, (b) a major fluency error, (c) a minor accuracy error, or (d) a minor fluency error. The formal definitions of these error types, along with the few-shot examples used for prompting, are provided in Appendix A.5. We then evaluate the resulting passages using MetricX-24 QE and Gemini-based AutoMQM. For simplicity, we report results for English–German and English–Chinese in Figures 3 and 4.
For the regression-based metric, we find that inserting a major error in the first segment tends to reduce score fluctuations as additional error-free segments are appended. In contrast, inserting a minor error has little influence on the existing length bias. This difference likely arises because major errors provide a strong signal that outweighs the effect of length bias by delineating a clearer boundary between erroneous and correct sentences, whereas minor errors contribute substantially weaker cues. Conversely, for the AutoMQM-based metric, the results suggest no consistent bias with respect to error severity. Finally, across both metrics, fluency errors appear to have a smaller impact on score changes than accuracy errors. A plausible explanation is that metrics are generally more sensitive to accuracy errors, while fluency errors are more subjective and tend to exhibit higher inter-rater disagreement in the training data.
Figure 6: Distribution of MetricX QE outputs and gold standard scores.
4.4 Bias in Translation Length
Experiment Setup.
In this section, we examine whether metrics exhibit biased preferences among translations of different lengths derived from the same source text. Specifically, we sample multiple hypothesis translations from each source as follows: we first pre-process the source and reference translations from the WMT24++ dataset into chunks that fit within a minimum context window of 500 tokens, discarding those that are too short (fewer than 200 tokens) to ensure meaningful variation while maintaining translation quality. For each source–reference pair, we prompt Gemini-2.5-Pro to generate five rephrased translations that preserve comparable accuracy and fluency, using temperature $t=1$ and top-$p=0.95$. Among these, we retain the two candidates with the minimum and maximum token lengths, thereby maximizing length difference while keeping translation quality comparable. All selected candidates are manually verified to ensure valid translations.
[…]
Results.
First, we note that in this experiment, the true quality value is set to $\theta=50\%$, indicating that an unbiased model should not distinguish between two translation variants of the same quality. However, across all evaluated settings, we observe a notable and consistent preference for shorter translations, with average scores ranging from 55% to 64%. This consistent deviation from the unbiased baseline suggests that length bias also exists with respect to translation length.
[…]
5 Analysis
Figure 7: Performance of MetricX-24 QE with and without length normalization. We report the slope of score changes to show that length normalization mitigates the magnitude of length bias.
5.1 Distribution of MQM Scores in WMT Data
To understand length bias of models for long-form inputs, we construct Blob-level MQM test sets using WMT data, each consisting of segments up to 500, 1k, or 2k tokens. Figure 6 presents the distribution of gold standard scores and MetricX-24 QE’s output scores on these test sets. We notice that (i) the distribution of output scores was skewed, and (ii) as the number of input tokens increases, the number of perfect translations (i.e., gold score = 0) in the source data diminishes, and the output distribution reflects this pattern, with the frequency of zero-score predictions dropping accordingly.
5.2 Length Normalization at Training
To validate our hypothesis that the bias stems from label density skew rather than architectural limitations, we apply a diagnostic intervention: simple length normalization. We posit that if a linear correction factor ($1/|h|$) is sufficient to remove the bias, the root cause is confirmed to be overfitting to the training data’s label distribution.
With normalization, the model is trained to predict the error density of a hypothesis translation, defined as $D(x,h)=R(x,h)/|h|$ where $R(x,h)$ is the error rating of hypothesis $h$ for source $x$ and $|h|$ denotes its segment length. At inference time, we rescale the prediction back to an error rating via $\hat{R}(x,h)=D(x,h)\cdot|h|$ to enable comparison with the unnormalized baseline. Figure 6 shows the gold score distributions with normalization, demonstrating greater consistency across test sets of varying lengths.
To further examine the effectiveness of normalization, we trained a MetricX QE model with the above training-time normalization. Figure 6 (Bottom-Left) demonstrates the impact of this intervention. Crucially, the normalized model’s score distribution mirrors the ground-truth human distribution (Figure 6, Bottom-Right): the curves for different lengths collapse onto a single, invariant density function. This confirms that length normalization does not merely dampen scores, but successfully recalibrates the model to match the fundamental statistical properties of human quality judgments—a property the unnormalized baseline fails to capture. By restoring this distributional alignment, the model learns to decouple error probability from sequence length.
Additionally, we replicate the experimental setup in Section 4.2 and report results in Figure 7. To highlight the effect of normalization, we compare the absolute average score difference across segment lengths (i.e., the slope of score changes). Results indicate that applying normalization reduces length bias beyond three segments, consistent with our earlier analysis of output score distributions.
6 Conclusion
In this work, we show that contemporary reference-free QE models are not length-invariant when assessing long-form translations: they inflate predicted error as translations grow longer and systematically favor shorter candidates for the same source. We trace this phenomenon to a training pipeline artifact: the under-representation of long, error-free examples in standard datasets. By demonstrating that a simple length-normalization step restores alignment with human ground-truth distributions, we provide a diagnostic proof that the bias is distributional rather than architectural. While standard aggregate benchmarks often miss these tail-end failures, our findings highlight the necessity of explicitly auditing length-invariance to ensure reliable evaluation in long-context applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Translation evaluation. Beyond technical improvements, our findings have potential implications for linguistic fairness and model safety. We demonstrate that current QE metrics systematically penalize length, which may result in the unfair devaluation of translations in languages that are naturally more verbose (e.g., German, Russian) compared to concise languages (e.g., English, Chinese). Furthermore, as QE metrics are increasingly employed as reward signals in Reinforcement Learning (RLHF), unmitigated length bias poses a risk of driving generation models toward excessive brevity, potentially leading to information loss or over-simplification of complex source concepts. By diagnosing and mitigating this bias, our work aims to ensure that automated evaluation remains equitable across diverse linguistic typologies and promotes reliability in automated training pipelines.
Appendix A Appendix
A.1 Length Bias in Reference-Based Metrics
When references are available, we compare reference-based, source-based (QE), and hybrid MT metrics under the experiment setup described in Section 4.2. Figure 8 shows that metrics with access to references exhibit attenuated length bias, plausibly because the reference translation provides an implicit prior on the expected length of a well-formed hypothesis. Nonetheless, these metrics continue to exhibit residual length bias, indicating that reference access alone does not fully eliminate the effect.
Figure 8: Performance of three classes of MetricX-24: reference-based, source-based (QE), and hybrid. We report the slope of score changes to illustrate that metrics with access to references exhibit reduced length bias, though the effect nonetheless persists.
A.2 Length Bias with respect to Model Size
Some of the QE models we study (e.g., CometKiwi and LLM-as-a-Judge) are only available in a single model size. In contrast, MetricX-24 provides multiple variants (Large, XL, and XXL). In this subsection, we examine how length bias varies with model capacity by repeating our main diagnostics using these MetricX-24 variants.
Experimental setup.
Unless otherwise stated, we use the bfloat16 variants of the Large and XL models for consistency with the publicly available bfloat16 XXL checkpoint. [^2] https://huggingface.co/google/metricx-24-hybrid-xxl-v2p6-bfloat16
Experiment 1: Bias in source/translation length.
We repeat the concatenation experiment from Section 3.2 using all three MetricX-24 variants. Table 1 reports the average predicted error as we concatenate five error-free segments. We observe a clear relationship between model size and the severity of the length-related penalty: smaller models exhibit stronger bias.
Model 1 2 3 4 5
Large 0.00 1.41 2.49 3.38 4.16
XL 0.00 1.02 1.61 2.24 2.61
XXL 0.00 0.55 0.80 1.26 1.47
Table 1: Model-size analysis on concatenating error-free segments (passage index 1155). Smaller MetricX-24 variants exhibit larger increases in predicted error as length grows.
Experiment 2: Bias in translation length preferences.
We repeat the reranking-style length preference experiment from Section 3.4. In an unbiased setting, a model should show roughly 50% preference for the shorter translation when two candidates are of comparable quality but differ in length. Table 2 reports the preference rates for the shorter translation at different length gaps. Consistent with Experiment 1, the smaller model exhibits a more pronounced length preference, while the XL and XXL variants behave similarly.
Model 2.5% 5% 7.5% 10% 12.5% 15%
Large 57.2 57.3 57.3 57.7 59.6 62.5
XL 53.5 53.6 53.4 54.4 55.9 56.5
XXL 54.3 54.0 53.4 54.0 56.8 57.8
Table 2: Model-size analysis on length preference: percentage preference for the shorter translation under different length differences. Values closer to 50% indicate less length bias.
Taken together, these results suggest that increasing model capacity can reduce length bias, but substantial bias remains even for the largest publicly available MetricX-24 variant, motivating model-agnostic mitigation strategies such as length-normalized training.
A.3 Proportion of Scores with Decreasing Trends in Data Filtering
Additionally, we analyze the proportion of documents exhibiting a decreasing trend in translation scores. For each language pair, we compare the score of passage 1 (containing segment 1) with passage 5 (containing segments 1–5), and report (i) the total number of documents and (ii) the proportion that display a decreasing trend. The results for MetricX-24 QE are summarized in Table 3. Overall, about 80% of the documents across 10 languages show a decreasing trend, indicating that length bias is pervasive in most cases.
Language # of Documents Proportion of scores with decreasing trends (%)
Aggregate 472 80.1
German (de_DE) 48 85.4
Chinese (zh_CN) 53 79.2
Spanish (es_MX) 47 74.5
French (fr_FR) 45 82.2
Japanese (ja_JP) 54 90.7
Korean (ko_KR) 46 82.6
Hindi (hi_IN) 37 75.7
Arabic (ar_EG) 48 68.8
Russian (ru_RU) 48 79.2
Portuguese (pt_BR) 46 80.4
Table 3: Proportion of scores with decreasing trends with MetricX-24 QE
A.4 Numerical Results for the RLHF and Reranking Experiments
Tables 4, 5, 6, 7 present the numerical results for the RLHF and Reranking experiments in Section 4.4 in the tables below. Numbers in brackets indicate the total number of instances.
Target Language Preference for Shorter Translations (%) by Length Difference
2.5% 5% 7.5% 10% 12.5% 15%
German (de_DE) 55.4 (101) 53.8 (93) 55.4 (74) 58.3 (48) 48.6 (35) 45.8 (24)
Chinese (zh_CN) 52.3 (86) 54.4 (79) 53.1 (64) 53.1 (49) 58.3 (36) 56.0 (25)
Spanish (es_MX) 49.0 (104) 51.2 (80) 59.1 (44) 60.9 (23) 66.7 (15) 63.6 (11)
French (fr_FR) 68.9 (106) 68.0 (97) 68.4 (76) 70.6 (51) 68.8 (32) 70.0 (20)
Japanese (ja_JP) 63.0 (81) 63.2 (76) 66.7 (54) 67.5 (40) 72.4 (29) 75.0 (20)
Korean (ko_KR) 54.3 (105) 55.1 (98) 56.8 (74) 48.0 (50) 45.5 (33) 50.0 (24)
Hindi (hi_IN) 44.8 (116) 45.5 (112) 47.5 (101) 48.3 (87) 48.7 (78) 44.8 (67)
Arabic (ar_EG) 70.9 (103) 72.0 (100) 71.1 (97) 70.8 (89) 69.1 (68) 67.9 (53)
Russian (ru_RU) 61.8 (102) 61.1 (95) 61.1 (72) 67.3 (55) 66.7 (36) 64.3 (28)
Portuguese (pt_PT) 46.2 (104) 47.4 (95) 49.3 (69) 50.0 (42) 61.9 (21) 61.1 (18)
Table 4: Results on MetricX-24 QE’s translation length preferences for EN $\rightarrow$ $X$ language pairs
Source Language Preference for Shorter Translations (%) by Length Difference
2.5% 5% 7.5% 10% 12.5% 15%
German (de_DE) 45.5 (101) 45.2 (93) 42.9 (63) 43.8 (48) 36.4 (33) 40.0 (25)
Chinese (zh_CN) 36.1 (83) 35.4 (82) 34.2 (73) 32.8 (61) 30.8 (52) 31.7 (41)
Spanish (es_MX) 63.1 (103) 62.6 (91) 62.9 (70) 68.4 (57) 62.2 (45) 65.8 (38)
French (fr_FR) 42.1 (107) 43.4 (99) 46.2 (78) 52.6 (57) 52.9 (34) 52.0 (25)
Japanese (ja_JP) 54.2 (83) 52.5 (80) 51.5 (68) 51.9 (52) 54.3 (35) 52.0 (25)
Korean (ko_KR) 57.1 (105) 60.2 (98) 60.2 (83) 59.7 (67) 61.8 (55) 65.2 (46)
Hindi (hi_IN) 75.9 (116) 75.7 (115) 77.5 (111) 78.5 (107) 77.9 (104) 78.4 (102)
Arabic (ar_EG) 59.2 (103) 61.2 (98) 61.7 (94) 65.5 (87) 68.8 (80) 70.1 (67)
Russian (ru_RU) 56.9 (102) 58.6 (99) 60.5 (86) 60.9 (69) 67.9 (56) 69.6 (46)
Portuguese (pt_PT) 59.0 (105) 58.8 (97) 53.2 (77) 51.6 (62) 53.1 (49) 58.5 (41)
Table 5: Results on MetricX-24 QE’s translation length preferences for $X$ $\rightarrow$ EN language pairs
Target Language Preference for Shorter Translations (%) by Length Difference
2.5% 5% 7.5% 10% 12.5% 15%
German (de_DE) 61.4 (101) 62.4 (93) 63.5 (74) 64.6 (48) 68.6 (35) 75.0 (24)
Chinese (zh_CN) 60.5 (86) 59.5 (79) 59.4 (64) 57.1 (49) 52.8 (36) 52.0 (25)
Spanish (es_MX) 53.8 (104) 58.8 (80) 68.2 (44) 65.2 (23) 66.7 (15) 63.6 (11)
French (fr_FR) 53.8 (106) 54.6 (97) 61.8 (76) 64.7 (51) 68.8 (32) 70.0 (20)
Japanese (ja_JP) 64.2 (81) 64.5 (76) 72.2 (54) 67.5 (40) 65.5 (29) 65.0 (20)
Korean (ko_KR) 57.1 (105) 57.1 (98) 54.1 (74) 62.0 (50) 69.7 (33) 66.7 (24)
Hindi (hi_IN) 60.3 (116) 61.6 (112) 60.4 (101) 58.6 (87) 61.5 (78) 59.7 (67)
Arabic (ar_EG) 49.5 (103) 49.0 (100) 50.5 (97) 50.6 (89) 48.5 (68) 49.1 (53)
Russian (ru_RU) 61.8 (102) 62.1 (95) 62.5 (72) 63.6 (55) 63.9 (36) 60.7 (28)
Portuguese (pt_PT) 59.6 (104) 60.0 (95) 58.0 (69) 57.1 (42) 47.6 (21) 55.6 (18)
Table 6: Results on CometKiwi’s translation length preferences for EN $\rightarrow$ $X$ language pairs
Source Language Preference for Shorter Translations (%) by Length Difference
2.5% 5% 7.5% 10% 12.5% 15%
German (de_DE) 68.3 (101) 69.9 (93) 77.8 (63) 75.0 (48) 69.7 (33) 64.0 (25)
Chinese (zh_CN) 74.7 (83) 74.4 (82) 72.6 (73) 75.4 (61) 75.0 (52) 75.6 (41)
Spanish (es_MX) 58.3 (103) 61.5 (91) 61.4 (70) 56.1 (57) 55.6 (45) 50.0 (38)
French (fr_FR) 73.8 (107) 72.7 (99) 70.5 (78) 75.4 (57) 70.6 (34) 60.0 (25)
Japanese (ja_JP) 77.1 (83) 77.5 (80) 77.9 (68) 76.9 (52) 77.1 (35) 80.0 (25)
Korean (ko_KR) 61.9 (105) 59.2 (98) 57.8 (83) 52.2 (67) 43.6 (55) 39.1 (46)
Hindi (hi_IN) 42.2 (116) 41.7 (115) 39.6 (111) 37.4 (107) 35.6 (104) 35.3 (102)
Arabic (ar_EG) 57.3 (103) 56.1 (98) 55.3 (94) 55.2 (87) 55.0 (80) 52.2 (67)
Russian (ru_RU) 67.6 (102) 66.7 (99) 65.1 (86) 71.0 (69) 67.9 (56) 67.4 (46)
Portuguese (pt_PT) 59.0 (105) 59.8 (97) 53.2 (77) 53.2 (62) 46.9 (49) 41.5 (41)
Table 7: Results on CometKiwi’s translation length preferences for $X$ $\rightarrow$ EN language pairs
A.5 Error Type / Severity Perturbation Prompt
Figures 9, 10, 11, 12, 13 present the prompts used for error perturbation with different types and severity using Gemini-2.5-pro.
Figure 9: Error perturbation prompt
Figure 10: Major accuracy prompt
Figure 11: Major fluency prompt
Figure 12: Minor accuracy prompt
Figure 13: Minor fluency prompt
A.6 LLM-as-a-Judge Prompt
Figure 14 presents the prompts used by LLM-as-a-Judge with Gemini-2.5-pro.
Figure 14: LLM-as-a-Judge Prompt